OpenStudy (sriram):

A 5 m ladder is bent on a wall Its base is pulled at 2m/s. Find rate of change of height at base=4m

OpenStudy (anonymous):



OpenStudy (anonymous):

Use Pythagoras

OpenStudy (anonymous):

\[T^2 + h^2 = 25 \] \[2T*\frac{dT}{dt} = - 2h*\frac{dh}{dt}\]Now just put the Values